OpenStudy (arianna842):

If the sum of the interior angle measures of a polygon is 3600°, how many sides does the polygon have? A. 22 sides B. 20 sides C. 18 sides D. 10 sides

OpenStudy (campbell_st):

well the formula fot the sum of the interior angles is \[Angle ~~Sum = (n - 2) \times 180\] where n is the number of sides substituting your information you get \[3600 = (n - 2) \times 180\] now just solve for n

OpenStudy (arianna842):

Thxs :P